OpenStudy (anonymous):

Which expressions would complete this equation so that the equation has no solution? 6(4x – 3) – 10x = Choose exactly two answers that are correct. A. 14x-15 B. 2(7x-10) C. 15x-14 D. 2(7x-9)

OpenStudy (luckycoins888):

6(4x – 3) – 10x = distribute the 6 first 6*4x = 24x 6 * (-3) =-18 24x-18-10x subtract the 10x from 24x 14x-18 since we know it's 14x-18 we can already see that answer A brings no solution. because 14x-18=14x-15 you can't find any x that will work. The lines are parallel to each other and will never cross answer C however is not correct because the lines will cross so you can find exactly 1 solution to it. so between B and D there's another correct answer 2(7x – 10) = 14x-20 these also never touch just like A so A and B both have no solutions the answers are A and B, since we are finding which ones DONT have solutions(:

OpenStudy (anonymous):

thanks foo i appreciate it

OpenStudy (luckycoins888):

Welcome (:

OpenStudy (anonymous):

lol

OpenStudy (luckycoins888):

xD